if A,B and C are mutually exclusive events and p(A) =0.2,p(B)=0.3 p(C)=0.2find (a) p(A union B union C),(b) p(Acomplement intersect (B union C)), (c) p(B union C complement ) complement


Answer by greenestamps(13327)



(a) The events are mutually exclusive, so P(A union B union C) is just P(A)+P(B)+P(C) = 0.2+0.3+0.2 = 0.7

(b) Again since the events are mutually exclusive, (A complement) intersect (B union C) is just B union C; so P[(A complement) intersect (B union C)] is P(B union C) = P(B)+P(C) = 0.3+0.2 = 0.5

(c) Once more, since the events are mutually exclusive, B union C complement is just C complement; the complement of that set is C. So P((B union C complement)complement) = P(C) = 0.2